# Vision Transformers in Precision Agriculture: A Comprehensive Survey

### v1

@@ Abstract
Detecting plant diseases is a crucial aspect of modern agriculture—it plays a key role in maintaining crop health and increasing overall yield. Traditional approaches, though still valuable, often rely on manual inspection or conventional machine learning techniques, both of which face limitations in scalability and accuracy. Recently, Vision Transformers (ViTs) have emerged as a promising alternative, offering benefits such as improved handling of long-range dependencies and better scalability for visual tasks. This survey explores the application of ViTs in precision agriculture, covering tasks from classification to detection and segmentation. We begin by introducing the foundational architecture of ViTs and discuss their transition from Natural Language Processing (NLP) to computer vision. The discussion includes the concept of inductive bias in traditional models like Convolutional Neural Networks (CNNs), and how ViTs mitigate these biases. We provide a comprehensive review of recent literature, focusing on key methodologies, datasets, and performance metrics. The survey also includes a comparative analysis of CNNs and ViT, with a look at hybrid models and performance enhancements. Technical challenges—such as data requirements, computational demands, and model interpretability—are addressed alongside potential solutions. Finally, we outline potential research directions and technological advancements that could further support the integration of ViTs in real-world agricultural settings. Our goal with this study is to offer practitioners and researchers a deeper understanding of how ViTs are poised to transform smart and precision agriculture.
@@ 1 Introduction
Plant diseases pose a significant challenge to global agriculture, affecting crop yields, food security, and the economic well-being of farmers. Early and accurate detection is essential to effectively manage these diseases and minimize their impact [11]. Traditionally, diagnosing plant diseases has relied on labor-intensive manual inspections by experts or conventional machine learning techniques—both of which come with limitations. Manual inspection is not only time-consuming but also subject to human error and inconsistency, making it impractical for large-scale agricultural operations [48, 23]. To address these limitations, researchers are increasingly turning to automated solutions that harness recent advances in machine learning and computer vision. CNN, a widely used traditional deep learning method, have shown some success in automating the disease detection process [77, 43]. However, these models often struggle with scalability and with capturing complex spatial patterns in images—both of which are vital for accurate detection and classification of plant diseases across varied environmental conditions [50]. This has created a growing demand for more robust, scalable, and intelligent systems for plant disease detection. In response, several studies have explored the potential of Deep Learning (DL) to meet these challenges. Some existing surveys provide general overviews of DL approaches, highlighting their advantages over traditional methods and outlining the progress made in applying these techniques to plant disease detection tasks. For example, one study [73] reviews how DL and Machine Learning (ML) techniques have been employed for plant disease detection from image data. The study describes the transition from handcrafted features with the traditional DL methods to deep models like CNNs, DBNs, and segmentation models. It also discusses the implementation of object detection algorithms like YOLO and Faster R-CNN, and segmentation architectures like U-Net and Mask R-CNN, for pixel-level accuracy-based detection quality improvement. Another work [34] discusses the application of ML and DL techniques in the early detection and classification of plant diseases, highlighting their importance in improving agricultural yield and reducing crop loss. It presents an extensive list of ML techniques such as Naïve Bayes, k-Nearest Neighbors, Decision Trees, Support Vector Machines, Random Forests, and Multi-layer Perceptrons, and explores their applications in plant disease classification using environmental and image inputs. The study also discusses preprocessing techniques and image segmentation methods, demonstrating how ML has enabled the automation of disease diagnosis and soil testing from models trained with crop data. The contrast between classical ML and DL methods shows that DL outperforms ML in terms of accuracy and scalability. It emphasizes the importance of addressing current deficiencies and proposes DL as a promising direction for creating automatic, systematic, and scalable plant disease diagnostic systems that support both scientists and farmers in agricultural decision-making. Another overview of studies [2] explores the use of DL techniques for plant disease detection and management in agriculture. It addresses key challenges including dataset quality, imaging sensors, model generalizability, estimation of severity, and performance comparison with human capabilities. Commonly used datasets are categorized, along with their limitations in practical scenarios. The study outlines various DL techniques—particularly CNN-based models—used for image classification, object detection, and semantic segmentation. It identifies shortcomings in existing models’ ability to generalize to field conditions due to overfitting and imbalanced data. Additionally, significant research areas are highlighted, such as universal severity estimation, multi-disease detection, and model generalization across diverse conditions. A separate study [44] explores the application of DL in plant disease recognition and classification, emphasizing its critical role in agricultural productivity and food security. It illustrates how DL improves upon traditional image processing techniques through automated feature extraction and enhanced objectivity and efficiency. The paper discusses foundational concepts, model performance criteria, and the evolution of DL in plant disease recognition. It also addresses the need for data augmentation to manage limited datasets and the use of visualization methods to improve model interpretability. The study concludes with future challenges, such as the need for more varied datasets, enhanced model robustness, and the integration of hyperspectral imaging for early detection.
@@
The use of ViTs in agriculture is a novel and rapidly emerging research area. To the best of our knowledge, no review article specifically focused on ViTs and their applications in precision agriculture currently exists, although some works briefly mention ViT. For example, a review [54] categorizes DL techniques applied to plant disease detection, organizing the literature into three tasks: classification, detection, and segmentation of plant diseases from leaf images. The review evaluates a broad range of DL models, with a strong focus on CNNs such as ResNet, VGG, and EfficientNet, alongside growing interest in ViTs. Another study [85] examines the use of DL and computer vision in plant disease detection, systematically analyzing the transition from traditional image processing to modern DL-based approaches. It assesses the suitability and performance of image processing techniques like thresholding, edge detection, and region-based segmentation in real-world agriculture. The study categorizes various DL architectures—including CNNs, Generative Adversarial Network (GAN), ViTs, and vision-language models—and highlights their scalability and practical implementation. It covers advanced models like YOLO, SSD, and Faster R-CNN for real-time object detection, and discusses the emerging relevance of ViTs for capturing global context in image analysis. However, while ViTs are introduced, the review lacks a deep evaluation of specific architectures or performance benchmarks in agricultural contexts.
In this work, we examine the use of ViT in the agriculture domain with the objective of identifying how this architecture has been used and adapted for agricultural applications such as crop classification, disease detection, yield prediction, and precision agriculture. With ViT’s success in computer vision across numerous domains, its application to transforming agricultural image analysis holds much promise. From reading these papers, we aspire to extract existing trends, model performance, and discover challenges particular to agricultural environments and datasets. This initial exploration will lay the groundwork for future studies and build a sound body of knowledge within this emerging interface of DL and agriculture.
@@
  • We have comprehensively addressed the concept of inductive biases and their role in machine learning models, especially CNNs and ViTs.
@@
The structure of this paper is organized as follows: Section 2 provides an overview of the Transformer architecture, followed by a discussion on ViTs. Section 3 explores the concept of inductive biases in machine learning, with a particular focus on how these biases manifest in CNNs and ViTs. In Section 4, we review the applications of ViTs in precision agriculture, categorizing the models into two groups: pure and hybrid models. Section 5 highlights key findings and outlines the open challenges identified in the reviewed literature. Finally, Section 6 presents the conclusions of the paper.
@@ 2.1 Transformer
The Transformer model, proposed by [86], brought a groundbreaking shift to NLP. Its core innovation is the self-attention mechanism, which enables the model to assess the significance of various words within a sentence during sequence encoding and generation. This mechanism addresses the limitations of earlier architectures, such as Recurrent Neural Networks (RNNs) and CNNs, which often struggled with long-range dependencies and parallelization challenges. By processing all tokens in a sequence simultaneously, self-attention makes the Transformer highly efficient and scalable. In the following subsection, we examine its main components.
@@ 2.1.3 Multi-Head Attention
    Parallel Attention Heads: pply the scaled dot-product attention mechanism independently to each set of queries, keys, and values in parallel. This allows the model to attend to different parts of the input sequence from different subspaces.
@@
    Concatenation: oncatenate the outputs of all $h$ attention heads, forming a single vector of combined information..
@@ 2.2 Vision Transformer
Inspired by the success of Transformers in NLP, [19] adapted the Transformer architecture for computer vision tasks, leading to the creation of ViT. Before this breakthrough, visual attention mechanisms were typically combined with convolutional networks or used to enhance specific components, while retaining the overall CNN structure. Early attention-based architectures had shown promise in computer vision [33], but their implementation on hardware accelerators often required complex and specialized engineering, as these systems were primarily optimized for convolutional operations. The key insight of ViT was to show that convolutional networks are not essential: a simple Transformer applied directly to sequences of image patches can achieve strong results in image classification. In ViT, images are divided into fixed-size patches, each treated as a token, similar to words in NLP. These patches are linearly embedded, combined with positional encodings, and processed through multiple Self-Attention layers. This design enables ViTs to effectively capture long-range dependencies and contextual relationships within images, leading to performance that often surpasses traditional CNNs on various classification benchmarks. By leveraging the Self-Attention mechanism of Transformers, ViTs have revolutionized the field of computer vision. The following section delves into the architecture of ViTs, explaining their key components and operations in detail. Figure 4 shows the structure of ViT. The architecture of ViT consists of the following key components: Patch Embedding, Positional Embedding, Transformer Encoder, and Classification Head.
@@ 2.2.1 Patch Embedding
Let denote $z^{P}_{0}\in\mathbb{R}^{D}$ the initial embedding of the $P^{th}$ image patch. It is obtained by flattening the patch and projecting it into a D-dimensional space using a learnable linear transformation:
@@ 3.1 Definition of Inductive Biases
Inductive biases refer to a set of built-in assumptions that a machine learning model creates, which can facilitate the generalization process, especially when working with limited data. We can consider these biases as prior knowledge that the model uses to make predictions about data that it hasn’t seen before. In general terms, inductive biases guide the learning algorithm to solutions that possess particular characteristics [7, 27].
@@ 3.2 Inductive Biases In CNNs and ViTs
In particular, two dominant inductive biases—locality and weight sharing—serve as architectural constraints that significantly reduce the number of learnable parameters and enhance the model’s generalization capabilities [20]. Locality is the design principle whereby each neuron in a convolutional layer only connects to a small, spatially localized subset of the input—the receptive field. This local computation enables CNNs to effectively extract low-level features such as edges, corners, and textures. As data passes through deeper layers, local features are hierarchically combined to detect more abstract and global representations [25]. Weight sharing, meanwhile, ensures the same filters (kernels) are employed at different spatial locations, which makes the model equivariant to translation. As a result, inductively learned patterns in one region of an image are recognizable wherever they are, thereby making the model stronger and more efficient, significantly reducing the number of trainable parameters, enabling it to generalize well, and improving computational efficiency [16]. Together, these inductive biases not only reduce the model’s complexity but also add prior knowledge about the structure of visual data, which accelerates training and improves performance, especially in scenarios with limited labeled data. Although CNNs’ intrinsic inductive biases can be very effective when data is limited. They can be limiting when there is sufficient data. In this situation, we need to choose the most effective inductive biases—not necessarily all of them [20]. The benefit of inductive biases tends to diminish with very large datasets, implying that transfer learning scenarios—where only a limited number of examples are available for the new distribution—provide a valuable context to assess the effectiveness and implementation of these biases [27]. Because of these characteristics, a randomly initialized CNN can perform well in object localization tasks without any learning [12]. Despite all these advantages, the locality characteristic of CNNs causes them to degrade in their ability to capture long-range dependencies [20].
@@ 3.3 Strategies to Mitigate the Lack of Inductive Biases in ViTs
ViTs lack the explicit inductive biases that are inherent in CNNs, such as locality and weight sharing, which can make it challenging for them to learn effectively, especially with limited data. One approach to mitigate this lack of inductive biases involves injecting convolutional inductive biases into ViTs. This can be done through convolutional patch embedding, which is a modification to how image data is turned into input tokens for the Transformer—replacing or enhancing the original flat patch embedding with convolutional layers to capture local spatial patterns more effectively—or by adding a convolutional stem at the beginning of the network to create hybrid models, which helps the model learn local features more effectively.
@@
Using hierarchical structures is another way to mitigate the lack of inductive biases by introducing a multi-stage architecture that processes visual information similarly to CNNs. Instead of attending globally from the start, these models apply self-attention within local windows and gradually merge or downsample tokens across layers, creating a spatial hierarchy [47]. This setup embeds locality and spatial coherence—key inductive biases of CNNs—allowing the model to first capture low-level features locally and then build up to more abstract, global representations [88]. Therefore, we can summarize the most common approaches that are in use for mitigating the lack of inductive biases in ViTs as:
  • Convolutional Patch Embedding: Replacing the simple patch embedding (linear projection) with convolutional layers to preserve local structures in the image, or using recursive token aggregation that behaves similarly to convolutions in how it captures local spatial relationships (convolutional-like patch embedding), is a method used in Tokens-to-Token ViT (T2T-ViT) [94].
  • Hybrid Models: As we mentioned before, CNNs are good at capturing local dependencies, and ViTs are good at capturing long-range global dependencies. Using a CNN or convolutional backbone to extract low-level features and then using this output as an input to a ViT is one of the ways of mitigating inductive biases in ViTs. The result of this is combining the spatial inductive biases of CNNs with the global modeling of ViTs. Adding a few convolutional inductive biases in the earlier steps of ViT’s processing can create a balance between inductive biases and the learning ability of Transformer blocks, and also change ViT’s optimization behavior [90].
@@
  • Hierarchical Structure: Designing ViTs with hierarchical architectures to process multi-scale features is what Swin Transformer: Hierarchical Vision Transformer using Shifted Windows (SwinT) does, which results in improved efficiency and local feature awareness [47].

### v2

@@ Abstract
Detecting plant diseases is a crucial aspect of modern agriculture, as it plays a key role in maintaining crop health and increasing overall yield. Traditional approaches, though still valuable, often rely on manual inspection or conventional machine learning techniques, both of which face limitations in scalability and accuracy. Recently, Vision Transformers (ViTs) have emerged as a promising alternative, offering advantages such as improved handling of long-range dependencies and better scalability for visual tasks. This review explores the application of ViTs in precision agriculture, covering a range of tasks. We begin by introducing the foundational architecture of ViTs and discussing their transition from Natural Language Processing (NLP) to Computer Vision. The discussion includes the concept of inductive bias in traditional models like Convolutional Neural Networks (CNNs), and how ViTs mitigate these biases. We provide a comprehensive review of recent literature, focusing on key methodologies, datasets, and performance metrics. This study also includes a comparative analysis of CNNs and ViTs, along with a review of hybrid models and performance enhancements. Technical challenges such as data requirements, computational demands, and model interpretability are addressed, along with potential solutions. Finally, we outline future research directions and technological advancements that could further support the integration of ViTs in real-world agricultural settings. Our goal with this study is to offer practitioners and researchers a deeper understanding of how ViTs are poised to transform smart and precision agriculture.
@@ 1 Introduction
Plant diseases pose a significant challenge to global agriculture, affecting crop yields, food security, and the economic well-being of farmers. Early and accurate detection is essential to effectively manage these diseases and minimize their impact [11]. Traditionally, diagnosing plant diseases has relied on labor-intensive manual inspections by experts or conventional machine learning techniques—both of which come with limitations. Manual inspection is not only time-consuming but also subject to human error and inconsistency, making it impractical for large-scale agricultural operations [48, 23]. To address these limitations, researchers are increasingly turning to automated solutions that harness recent advances in Machine Learning (ML) and computer vision. CNN, a widely used traditional Deep Learning (DL) method, have shown some success in automating the disease detection process [77, 43]. However, these models often struggle with scalability and with capturing complex spatial patterns in images—both of which are vital for accurate detection and classification of plant diseases across varied environmental conditions [50]. This has created a growing demand for more robust, scalable, and intelligent systems for plant disease detection. In response, several studies have explored the potential of DL to meet these challenges. Some existing surveys provide general overviews of DL approaches, highlighting their advantages over traditional methods and outlining the progress made in applying these techniques to plant disease detection tasks. For example, one study [73] reviews how DL and ML techniques have been employed for plant disease detection from image data. The study describes the transition from handcrafted features with the traditional DL methods to deep models like CNNs, DBNs, and segmentation models. It also discusses the implementation of object detection algorithms like YOLO and Faster R-CNN, and segmentation architectures like U-Net and Mask R-CNN, for pixel-level accuracy-based detection quality improvement. Another work [34] discusses the application of ML and DL techniques in the early detection and classification of plant diseases, highlighting their importance in improving agricultural yield and reducing crop loss. It presents an extensive list of ML techniques such as Naïve Bayes, k-Nearest Neighbors, Decision Trees, Support Vector Machines, Random Forests, and Multi-layer Perceptrons, and explores their applications in plant disease classification using environmental and image inputs. The study also discusses preprocessing techniques and image segmentation methods, demonstrating how ML has enabled the automation of disease diagnosis and soil testing from models trained with crop data. The contrast between classical ML and DL methods shows that DL outperforms ML in terms of accuracy and scalability. It emphasizes the importance of addressing current deficiencies and proposes DL as a promising direction for creating automatic, systematic, and scalable plant disease diagnostic systems that support both scientists and farmers in agricultural decision-making. Another overview of studies [2] explores the use of DL techniques for plant disease detection and management in agriculture. It addresses key challenges including dataset quality, imaging sensors, model generalizability, estimation of severity, and performance comparison with human capabilities. Commonly used datasets are categorized, along with their limitations in practical scenarios. The study outlines various DL techniques—particularly CNN-based models—used for image classification, object detection, and semantic segmentation. It identifies shortcomings in existing models’ ability to generalize to field conditions due to overfitting and imbalanced data. Additionally, significant research areas are highlighted, such as universal severity estimation, multi-disease detection, and model generalization across diverse conditions. A separate study [44] explores the application of DL in plant disease recognition and classification, emphasizing its critical role in agricultural productivity and food security. It illustrates how DL improves upon traditional image processing techniques through automated feature extraction and enhanced objectivity and efficiency. The paper discusses foundational concepts, model performance criteria, and the evolution of DL in plant disease recognition. It also addresses the need for data augmentation to manage limited datasets and the use of visualization methods to improve model interpretability. The study concludes with future challenges, such as the need for more varied datasets, enhanced model robustness, and the integration of hyperspectral imaging for early detection.
@@
The use of ViTs in agriculture is a novel and rapidly emerging research area. To the best of our knowledge, there are a few works whose focus is on ViTs; however, most of them mention it only briefly and do not provide a comprehensive analysis of their applications in precision agriculture. For example, a review [54] categorizes DL techniques applied to plant disease detection, organizing the literature into three tasks: classification, detection, and segmentation of plant diseases from leaf images. The review evaluates a broad range of DL models, with a strong focus on CNNs such as ResNet, VGG, and EfficientNet, alongside growing interest in ViTs. Another study [85] examines the use of DL and computer vision in plant disease detection, systematically analyzing the transition from traditional image processing to modern DL-based approaches. It assesses the suitability and performance of image processing techniques like thresholding, edge detection, and region-based segmentation in real-world agriculture. The study categorizes various DL architectures—including CNNs, Generative Adversarial Network (GAN), ViTs, and Vision Language Models—and highlights their scalability and practical implementation. It covers advanced models like YOLO and Faster R-CNN for real-time object detection, and discusses the emerging relevance of ViTs for capturing global context in image analysis. However, while ViTs are introduced, the review lacks a deep evaluation of specific architectures or performance benchmarks in agricultural contexts.
In this work, we examine the use of ViT in the agriculture domain with the objective of identifying how this architecture has been used and adapted for agricultural applications such as crop classification, disease detection, yield prediction, and precision agriculture. With ViT’s success in computer vision across numerous domains, its application to transforming agricultural image analysis holds much promise. By analyzing the existing literature, we aspire to extract current trends, evaluate model performance, and identify challenges specific to agricultural environments and datasets. This initial exploration will lay the groundwork for future studies and build a sound body of knowledge within this emerging interface of DL and agriculture.
@@
  • We have comprehensively addressed the concept of inductive biases and their role in ML models, especially CNNs and ViTs.
@@
The structure of this paper is organized as follows: Section 2 provides an overview of the Transformer architecture, followed by a discussion on ViTs. Section 3 explores the concept of inductive biases in ML, with a particular focus on how these biases manifest in CNNs and ViTs. In Section 4, we review the applications of ViTs in precision agriculture, categorizing the models into two groups: pure and hybrid models. Section 5 highlights key findings and outlines the open challenges identified in the reviewed literature. Finally, Section 6 presents the conclusions of the paper.
@@ 2.1 Transformer
The Transformer model, proposed by [86], brought a groundbreaking shift to NLP. Its core innovation is the self-attention mechanism, which enables the model to assess the significance of various words within a sentence during sequence encoding and generation. This mechanism addresses the limitations of earlier architectures, such as Recurrent Neural Networks (RNNs) and CNNs, which often struggled with long-range dependencies and parallelization challenges. By processing all tokens in a sequence simultaneously, self-attention enables the Transformer to model long-range dependencies more effectively, making the architecture highly efficient and scalable for various sequence tasks. In the following subsection, we examine its main components.
@@ 2.1.3 Multi-Head Attention
    Parallel Attention Heads: Apply the scaled dot-product attention mechanism independently to each set of queries, keys, and values in parallel. This allows the model to attend to different parts of the input sequence from different subspaces.
@@
    Concatenation: Concatenate the outputs of all $h$ attention heads, forming a single vector of combined information.
@@ 2.2 Vision Transformer
Inspired by the success of Transformers in NLP, [19] adapted the Transformer architecture for computer vision tasks, leading to the creation of ViT. Before this breakthrough, visual attention mechanisms were typically combined with convolutional networks or used to enhance specific components, while retaining the overall CNN structure. Early attention-based architectures had shown promise in computer vision [33], but their implementation on hardware accelerators often required complex and specialized engineering, as these systems were primarily optimized for convolutional operations. The key insight of ViT was to show that convolutional networks are not essential: a simple Transformer applied directly to sequences of image patches can achieve strong results in image classification. In ViT, images are divided into fixed-size patches, each treated as a token, similar to words in NLP. These patches are linearly embedded, combined with positional embeddings, and processed through multiple Self-Attention layers. This design enables ViTs to effectively capture long-range dependencies and contextual relationships within images, leading to performance that often surpasses traditional CNNs on various classification benchmarks. By leveraging the Self-Attention mechanism of Transformers, ViTs have revolutionized the field of computer vision. The following section delves into the architecture of ViTs, explaining their key components and operations in detail. Figure 4 shows the structure of ViT. The architecture of ViT consists of the following key components: Patch Embedding, Positional Embedding, Transformer Encoder, and Classification Head.
@@ 2.2.1 Patch Embedding
Let $z^{P}_{0}\in\mathbb{R}^{D}$ denote the initial embedding of the $P^{th}$ image patch. It is obtained by flattening the patch and projecting it into a D-dimensional space using a learnable linear transformation:
@@ 3.1 Definition of Inductive Biases
Inductive biases refer to a set of built-in assumptions that a ML model creates, which can facilitate the generalization process, especially when working with limited data. We can consider these biases as prior knowledge that the model uses to make predictions about data that it hasn’t seen before. In general terms, inductive biases guide the learning algorithm to solutions that possess particular characteristics [7, 27].
@@ 3.2 Inductive Biases In CNNs and ViTs
In particular, two dominant inductive biases—locality and weight sharing—serve as architectural constraints that significantly reduce the number of learnable parameters and enhance the model’s generalization capabilities [20]. Locality is the design principle whereby each neuron in a convolutional layer only connects to a small, spatially localized subset of the input—the receptive field. This local computation enables CNNs to effectively extract low-level features such as edges, corners, and textures. As data passes through deeper layers, local features are hierarchically combined to detect more abstract and global representations [25]. Weight sharing, meanwhile, ensures the same filters (kernels) are employed at different spatial locations, which makes the model equivariant to translation. As a result, inductively learned patterns in one region of an image are recognizable wherever they are, thereby making the model stronger and more efficient, significantly reducing the number of trainable parameters, enabling it to generalize well, and improving computational efficiency [16]. Together, these inductive biases not only reduce the model’s complexity but also add prior knowledge about the structure of visual data, which accelerates training and improves performance, especially in scenarios with limited labeled data. Although CNNs’ intrinsic inductive biases can be very effective when data is insufficient, they can be limiting when there is sufficient data. In this situation, we need to choose the most effective inductive biases—not necessarily all of them. [20]. The benefit of inductive biases tends to diminish with very large datasets, implying that transfer learning scenarios—where only a limited number of examples are available for the new distribution—provide a valuable context to assess the effectiveness and implementation of these biases [27]. It is due of these characteristics that a randomly initialized CNN can perform well in object localization tasks without any learning [12]. Despite all these advantages, the locality characteristic of CNNs causes them to degrade in their ability to capture long-range dependencies [20].
@@ 3.3 Strategies to Mitigate the Lack of Inductive Biases in ViTs
ViTs lack the explicit inductive biases that are inherent in CNNs, such as locality and weight sharing, which can make it challenging for them to learn effectively, especially with limited data. One approach to mitigate this lack of inductive biases involves injecting convolutional inductive biases into ViTs. This can be done through convolutional patch embedding, which modifies how image data is turned into input tokens for the Transformer. Specifically, it replaces or enhances the original flat patch embedding with convolutional layers to capture local spatial patterns more effectively. Alternatively, a convolutional stem can be added at the beginning of the network to create hybrid models, helping the model learn local features more effectively.
@@
Using hierarchical structures is another way to mitigate the lack of inductive biases by introducing a multi-stage architecture that processes visual information similarly to CNNs. Instead of attending globally from the start, these models apply self-attention within local windows and gradually merge or downsample tokens across layers, creating a spatial hierarchy [47]. This setup embeds locality and spatial coherence—key inductive biases of CNNs—allowing the model to first capture low-level features locally and then build up to more abstract, global representations [88]. Therefore, the most common approaches used to mitigate the lack of inductive biases in ViTs can be summarized as follows:
  • Convolutional Patch Embedding: It means replacing simple patch embedding with convolutional layers to preserve local structures in the image, or using recursive token aggregation—which behaves similarly to convolutions in capturing local spatial relationships—is a method employed in models like Tokens-to-Token ViT (T2T-ViT) [94].
  • Hybrid Models: As mentioned earlier, CNNs are effective at capturing local dependencies, while ViTs excel at modeling long-range global dependencies. One way to mitigate the lack of inductive biases in ViTs is to use a CNN or convolutional backbone to extract low-level features, then feed this output into a ViT. This approach combines the spatial inductive biases of CNNs with the global modeling capabilities of ViTs. Introducing a few convolutional inductive biases in the early stages of ViT processing can strike a balance between inductive bias and the learning capacity of Transformer blocks, while also influencing the optimization behavior of ViTs [90].
@@
  • Hierarchical Structure: Designing ViTs with hierarchical architectures to process multi-scale features is what Swin Transformer (SwinT) does, which results in improved efficiency and local feature awareness [47].